# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v2

@@ II.1 Nuclear equation of state phenomenology
Following previous investigations, we characterise the nuclear equation of state with a low-dimensional phenomenological model. Rather than employ a sophisticated but high-dimensional phenomenological approach based on microphysics-motivated calculations with many tunable parameters such as relativistic mean field theory [84, 85, 86, 87, 88, 89, 90, 91], for simplicity and lacking observationally-driven necessity, we adopt two distinct analytic representations.
@@
In this work we principally employ a specific and widely-used spectral parameterization [76], referred to as the $\Gamma$-spectral parameterization from here onwards, in which the natural log of the zero-temperature adiabatic index $\Gamma=d\ln p/d\ln n$ is expressed as a power series in $\log p$:
@@
For this model, we adopt a uniform prior in its hyperparameters, with limits consistent with previous studies [92, 2]. Though we do not explicitly enforce other customary constraints (e.g., on the adiabatic index), our TOV solver requires EoS that are nearly causal, so our calculations de facto reject highly unphysical samples as implausible. To illustrate the impact of model and prior systematics, we also employ another spectral formulation, referred to as $\Upsilon$-spectral from here onwards (though denoted the “improved causal spectral” formulation in Ref. [78]), designed to ensure smooth transition from the fixed low density piece to the fit higher density part and to ensure causality all throughout. In particular we use the pressure based formulation described in [78], where the spectral expansion is constructed for the sound speed dependent parameter $\Upsilon(p)$ defined as
$\Upsilon(p)=\frac{c^{2}-v^{2}(p)}{v^{2}(p)}~{}~{}~{}\,$ (2)
@@
This decomposition is very similar to the previous causal-$\Gamma$-spectral representation [77] with the only difference being that the zeroth coefficient, $\lambda_{0}$, is predetermined. Here $\lambda_{0}$ is fixed such that sound speed continuity is ensured at the matching point, $p_{0}$, between the low density and high density pieces
@@
We restrict our analyses to the first four independent parameter in this decomposition, $k=1,2,3,4$, in order to capture systematic features of the EoSs while keeping the exploration manageable. We choose a uniform prior for these four spectral parameters, over $\lambda_{1}\in[-3,3.2]$, $\lambda_{2}\in[-4,2]$, $\lambda_{3}\in[-0.2,0.8]$, and $\lambda_{4}\in[-0.05,0.02]$, where these ranges of each parameter are motivated by covering existing EOS as well as the $\Gamma$-spectral posterior for GW170817 [93]. Figure 1 shows our implied prior on $p(\rho)$ and nonrotating neutron star gravitational mass and radius. We use these two parametric approaches to qualitatively assess the kinds of systematic uncertainty exhibited among some of the commonly used parameterizations, such as piecewise polytropic [94], dynamical-piecewise polytropic [95, 92], and similar spectral methods [76, 77, 78] which adopt different basis functions or parameters [96].
@@
Bottom panel: Mass-Radius prior range for the EoSs in the above range. The grey shaded regions indicate various constraints on single-stable NS; ‘BH limit’ is where the Schwarzchild radius is larger than the NS radius, ‘Buchdahl limit’ is the range where an NS would need infinite pressure [97], and ‘Causality limit’ is obtained by finding the maximally compact (M/R) solution for Mass-Radius in the limiting physical range of pressure-density [98].
@@
We follow previous work that employs Gaussian quadrature to evaluate the energy-density at given pressure on the lalsuite implementation [2, 92]. We adopt a fixed low-density equation of state with the first four terms in the expansion determining the EoS. This parameterization has been previously constrained by comparison of its input GW170817 [2] and NICER measurements [37, 45]. For all EoSs, we solve the Tolman-Oppenheimer-Volkoff (TOV) equation [99, 100] for nonrotating neutron star structure in hydrostatic equilibrium using RePrimAnd [101, 102] to find one-parameter families of gravitational mass $M$, radius $R$, and tidal deformability $\Lambda$. The output for an equation of state $\alpha$ is characterized by a one-parameter family $R_{\alpha}(s),M_{\alpha}(s)$, where $s$ is a fiducial parameter characterizing our dimensionless prior. For simplicity, in this work we adopt the pseudo-enthalpy h, and only investigate the lowest stable branch of solutions.
What is a measurement of an equation of state?
Bayesian inference frameworks provide a likelihood and hence (with a prior) probability for each equation of state function. We can therefore interpret observations as providing information about a credible interval on the space of functions: for example, deducing a volume of EoS which contain 90% of all (joint) posterior probability. Figure 1 shows our 100 percentile prior EoS distribution. This interpretation, however, conflicts with the customary approach to rendering and reporting conclusions about the EoS, which are presented as marginal distributions for (uncorrelated) point measurements: for example, the distribution of pressure at a given density, or the radius of NS of a fiducial mass. As highlighted in previous work [5], the superposition of point measurements neglects important information including correlations between different densities (for $p(\rho)$) or masses (for $R(M)$). As a concrete example for any two densities $\rho_{i}$, the posterior pressures $\ln p(\rho_{1}),\ln p(\rho_{2})$ will be correlated by measurements, above and beyond any correlations introduced by our prior. Nonetheless, in this work to simplify comparison with previously-reported results, except for Figure 1 we render our conclusions using this point-measurement approach. Thus for example a confidence interval on $p(\rho)$ follows by taking all weighted EoS realizations $(p_{k}(\rho),w_{k})$ and, for each reference density $\rho_{i}$, constructing a confidence interval for $p(\rho_{i})$ using the weighted sample at that density. Similarly but not identical are confidence intervals on $\rho(p)$, constructed similarly but using the inverse relationship.
The conclusions about the EOS posterior described in the preceding paragraph assume only a prior on the functional space of EoS. However, many discussions of the NS mass-radius distribution adopt priors along the family of stable NS by, for example, adopting a prior on the NS central density; see, e.g., [103, 74, 45]. As discussed below, these analyses are in effect adopting a prior on the NS mass distribution. As a result, in order to replicate their results, we must present our results in a comparable fashion. Specifically, rather than construct a 90% credible interval for the NS radius at each fixed mass, we must instead construct a two-dimensional marginal distribution in both mass and radius. This two-dimensional posterior distribution will explicitly depend on the choice of NS mass distribution (or equivalently NS central density distribution).
II.2 Neutron star population models
Deducing the equation of state from observations of neutron star properties always requires some model for the neutron star population: empirical observables limit but do not uniquely determine the neutron star mass, spin, radius, and tidal deformability. In this work, unless otherwise noted, we will assume the isolated (and binary) neutron star population is uniformly distributed in gravitational mass $m$ between fixed, EoS-independent limits $m\in[m_{min},m_{max}]$. We specifically adopt $m_{min}=0.4~{}{\rm M_{\odot}}$ and $m_{max}=2.1~{}{\rm M_{\odot}}$, with the lower and upper mass limits chosen because of the collection of NSs we will use to constrain the nuclear EoS. Similarly, unless otherwise noted, we will assume the (aligned component) of neutron star spin is distributed isotropically in a sphere and uniformly in magnitude, with a maximum dimensionless value $a=S^{2}/m<0.05$. This fiducial population closely resembles the prior assumptions used when inferring the nature of low-mass gravitational wave sources. We emphasize that our prior represents our assumptions about the astrophysical distribution of neutron stars as realized in nature, which need not reach the maximum mass allowed by the nuclear equation of state.
For context, other groups have adopted a wide range of different implicit or explicit priors for the neutron star mass distribution. For example, some adopt priors on the neutron star central density, not the neutron star mass (e.g., a quadratic prior [74, 45] or a log-uniform prior [103] between the maximum central density and some reference density). Alternatively, even within the context of nominally uniform mass distributions, others [104] have specifically assumed $p(m|{\cal Y})$ is uniform between a fixed minimum mass and the EoS-dependent maximum mass $M_{max}({\cal Y})$, where ${\cal Y}$ here and henceforth denotes the EoS or its parameterization. In other words, this model asserts all NSs that can exist under an EoS do exist. In terms of the unit step function $\Theta_{I}(x)=1$ for $x\in I$ ($I$ being the interval) and 0 otherwise, their one- and two-dimensional priors are $p(m)=\Theta_{I}(m)/|I|$ and $p(m_{1},m_{2})=2\Theta_{I}(m_{1})\Theta_{I}(m_{2})\Theta(m_{1}-m_{2})/|I|^{2}$, where $I=[m_{min},M_{max}({\cal Y})]$ is the interval between $m_{min}$ and the EoS-dependent maximum mass. Compared to our fiducial approach, both of these aggressive population models introduce EoS-dependent normalization factors, which weakly penalizes models with large $M_{\rm max}({\cal Y})$ for single observations and strongly penalize them for many observations. Additionally, these aggressive population models should be strongly constrained by the absence of massive neutron stars; see, e.g., Alsing et al. [105, 106] and references therein. For broader context regarding prior choices, Wysocki et al. [93] demonstrated that, if not known a priori, the neutron star mass distribution and EoS must be jointly inferred, to avoid biasing the deduced EoS. In the main text and in appendices we will briefly discuss how alternative mass priors can impacts our results.
Several sources of external information provide good reason to adopt alternative informed priors. For example, several authors have proposed adopting comparable masses for binary neutron stars, motivated by galactic observations and conventional field formation scenarios which strongly favor nearly equal masses [107, 108, 109, 110, 111, 107, 108]. Similar arguments have been employed to rule out large or preferentially negatively-oriented spins [112]. (Of course, the limited statistics available from the GW census are consistent with a much broader mass ratio distribution [79, 113] and provide minimal constraints on plausible NS spin orientations; additionally, non-merging galactic BNS binaries have a much broader mass ratio distribution [114].) At the other extreme, because the kilonova AT2017gfo associated with GW170817 suggests very massive dynamical ejecta and since most predictions for this ejecta are suppressed for $q\simeq 1$, others have suggested favoring asymmetric masses where the dynamical ejecta isn’t suppressed [115, 116]. On similar grounds based on AT2017gfo, others have demonstrated that a strongly-spinning secondary neutron star ($m_{2}$) will produce substantially enhanced ejecta, as its outer layers are less bound [117, 118]. Strongly spinning NS mergers can occur for example if the secondary does not have time to spin down, due to fortuitous NS kicks [119] or Kozai-Lidov effects in hierarchical triples [120]. (Such a rapid formation scenario isn’t consistent with the large geometric offset for AT2017gfo but could occur in general; as noted in [120, 121] this rapid formation scenario would reconcile NS merger formation with r-process-enriched but metal-poor stars.) Particularly given the systematics associated with contemporary eject calculations, an exhaustive quantitative investigation of these additional speculative sources of prior information merits its own dedicated study.
Motivated by the discussion in the preceding paragraphs, we will analyze GW170817 with combinations of several alternative priors. To focus on high mass ratios only, we will perform some analyses with a mass ratio distribution $p(q)$ for AT2017gfo that is consistent with a uniform distribution of component masses but excludes $m_{2}/m_{1}>0.6$. (At GW170817’s source-frame chirp mass ${\cal M}_{c}=1.186~{}{\rm M_{\odot}}$, this constraint requires more extreme choices for source-frame primary and secondary mass: $m_{2}<1.062~{}{\rm M_{\odot}}$ and $m_{1}>1.77~{}{\rm M_{\odot}}$.) Conversely, when requiring GW170817’s mass ratio prior distribution be constrained to the range expected for GW-detected BNS sources derived from the galactic population, we require $m_{2}/m_{1}>0.9$ [107]. To allow for extreme spins, we consider an isotropic uniform spin magnitude prior which extends to $0.4$. To strongly disfavor extreme spin-orbit misalignment, we consider $\chi_{i,z}>0$. Operationally, all of these astrophysical prior choices can be implemented as alternative options for the upper and lower bounds adopted for $q,\chi_{i,z}$, when analyzing GW170817 with a nonprecessing GW model.
II.3 Neutron star mass-radius constraints
NICER observations of periodic X-ray emission on millisecond pulsars provides information about their masses and radii. Though their data product is reported as a weighted posterior distribution, we find these results can be well approximated by a simple Gaussian distribution, as illustrated in Figure 2 for the pulsar PSR-J0030 (red). In Appendix A we describe the best-fitting gaussian to each NICER (alongside HESS) observation, and the specific method we used to deduce its parameters. In terms of this Gaussian likelihood $L(M,R)$, we can evaluate the marginal likelihood for each equation of state $\alpha$ according to [122, 123, 63]
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\delta(M-M_{\alpha}(s))\mbox{d}M\mbox{d}R\bigg{]}p(s)\mbox{d}s$
$=\int\mathcal{L}(M(s),R_{\alpha}(M(s)))~{}~{}p(s)\frac{\mbox{d}M}{\mbox{d}M_{\alpha}/\mbox{d}s}$
$={\cal F}\int\mathcal{L}(M(s),R_{\alpha}(M(s)))~{}~{}\mbox{d}M$ (6)
where $s$ is a fiducial parameter characterizing our dimensionless prior along each one-parameter family of each EoS; $p(s)$ is a prior along that one-parameter family; and in the last line we’ve chosen $p(s)$ to be compatible with our uniform prior over the neutron star mass, with ${\cal F}$ the appropriate normalization constant in mass space: ${\cal F}=1/(m_{max}-m_{min})$. For simplicity, in this work we adopt the gravitational mass $M$ as our fiducial prior (i.e., so the denominator $(dM_{\alpha}/ds)/p(s)$ is constant), while recognizing the utility of other global choices like the pseudo-enthalpy. This choice of prior along each family strongly downweights contributions to the evidence near turning points in the mass-radius plane at the high and low mass ends of the curve(i.e., $dM/dR\simeq 0$). That said, our choice of prior enables us to use the same functional form for all equations of state, independent of their maximum mass.
For comparison, other priors will inherit different priori mass distributions from their prior on the assumed parameterization $s$. If instead one follows Ref. [74] and chooses $s$ to be the central density and $p(\rho_{c})$ to be some prior on that central density, then the implied mass distribution is proportional to $p(\rho_{c})/dM/d\rho_{c}$. Equivalently, the cumulative distribution function for any individual equation of state satisfies $P(<\rho_{c})=P(<M(\rho_{c}))$, for simplicity assuming a single-branched relationship between mass and central density. The specific prior adopted in [74] very strongly favors massive binaries, where $dR/dM$ is large.
Most neutron star mass measurements provide information only about the empirical pulsar star mass distribution, rather than any direct insight into the nuclear equation of state. However, a few very massive pulsars’ existence provides useful insight into the neutron star maximum mass, bounding it from below: J0740 [124, 125], with pulsar mass consistent with a Gaussian with mean and standard deviation $2.14~{}{\rm M_{\odot}}$ and $0.1~{}{\rm M_{\odot}}$; J0348+0432 [126], with pulsar mass $2.01~{}{\rm M_{\odot}}\pm 0.04~{}{\rm M_{\odot}}$; and J1614 [127], with pulsar mass $1.908~{}{\rm M_{\odot}}\pm 0.016~{}{\rm M_{\odot}}$. We follow [28] to express the likelihood of an EoS consistent with these observations as the product of standard normal distributions:
where $\Phi(z)$ is the cumulative distribution of the standard normal distribution for each maximum mass measurement and are shown in Fig. 2 as horizontal bands.
II.4 Gravitational wave observations
Gravitational wave observations of GW170817 constrain the binary masses $m_{i}$ and tidal deformabilities $\Lambda_{i}$ for each component of the binary [1, 2, 3, 128]. Specifically, the presence of matter at leading order impacts the binary’s inspiral, through each components’ dimensionless tidal deformability parameter $\Lambda=(2/3)k_{2}(c^{2}R/GM)^{5}$ [129], where $k_{2}$ is the $l=2$ Love number, $M$ is the mass, and $R$ is the radius. The leading order contribution to the phase evolution of an GW inspiral is given by the weighted combination of $\Lambda$ terms
The impact of tidal effects are included in many conventional approximate phenomenological estimates of outgoing radiation from merging compact binaries. In this work, we will focus on analyses with a contemporary state-of-the-art model, IMRPhenomPv2_NRTidalv2 [130, 131], which incorporates precession physics but omits higher-order modes. We note other contemporary models like NRHybSur3dq8Tidal [132] and TEOBResumS [133] include these higher-order modes and tidal effects.
Using the marginal likelihood $L(X)$ obtained by interpolating the RIFT marginal likelihoods and the EoS-derived relationship connecting $\Lambda(m|{\cal Y})$ to the neutron star gravitational mass $m$ and equation of state model ${\cal Y}$, we compute a marginal evidence

### v3

@@ 2.7 Nuclear equation of state phenomenology
Following previous investigations, when parameterizing a nuclear EOS $\alpha$, we characterise it with a low-dimensional phenomenological model. Rather than employ a sophisticated but high-dimensional phenomenological approach based on microphysics-motivated calculations with many tunable parameters such as relativistic mean field theory (1989RPPh...52..439R; 2003RvMP...75..121B; 2022PhRvC.106e5804A; 2023PhRvC.108b5809Z; 2023ApJ...943..163Z; 2018PhRvC..97b5805M; 2024MNRAS.529.4650H; 2023PhRvD.107j3018M), for simplicity and lacking observationally-driven necessity, we adopt two distinct analytic representations.
@@
In this work we principally employ a specific and widely-used spectral parameterization (Lindblom:2010bb), referred to as the $\Gamma$-spectral parameterization from here onwards, in which the natural log of the zero-temperature adiabatic index $\Gamma=d\ln p/d\ln n$ is expressed as a power series in $\log p$:
@@
For this model, we adopt a uniform prior in its hyperparameters, with limits consistent with previous studies (Carney:2018sdv; LIGO-GW170817-SourceProperties). Though we do not explicitly enforce other customary constraints (e.g., on the adiabatic index), our TOV solver requires EoS that are nearly causal, so our calculations de facto reject highly unphysical samples as implausible. As noted in Appendix B of2020arXiv200101747W, this EoS model family and prior has a narrow, strongly correlated subspace of physical realizations: only a fraction $5\times 10^{-5}$ of the prior volume is physical. For this reason, adaptive or well-tuned sampling will be an essential to obtain robust results with this framework. To illustrate the impact of model and prior systematics, we also employ another spectral formulation, referred to as $\Upsilon$-spectral from here onwards (though denoted the “improved causal spectral” formulation in PhysRevD.105.063031), designed to ensure smooth transition from the fixed low density piece to the fit higher density part and to ensure causality all throughout. In particular we use the pressure based formulation described in (PhysRevD.105.063031), where the spectral expansion is constructed for the sound speed, $v(p)$, dependent parameter $\Upsilon(p)$ defined as
$\Upsilon(p)=\frac{c^{2}-v^{2}(p)}{v^{2}(p)}\penalty 10000\ \penalty 10000\ \penalty 10000\ \,$ (9)
@@
This decomposition is very similar to the previous causal-$\Gamma$-spectral representation (Lindblom:2018rfr) with the only difference being that the zeroth coefficient, $\lambda_{0}$, is predetermined. Here $\lambda_{0}$ is fixed such that sound speed continuity is ensured at the matching point, $p_{0}$, between the low density and high density pieces
@@
We restrict our analyses to the first four independent parameter in this decomposition, $k=1,2,3,4$, in order to capture systematic features of the EoSs while keeping the exploration manageable. For simplicity, we choose a uniform prior for these four spectral parameters, over $\lambda_{1}\in[-3,3.2]$, $\lambda_{2}\in[-4,2]$, $\lambda_{3}\in[-0.2,0.8]$, and $\lambda_{4}\in[-0.05,0.02]$, where these ranges of each parameter are motivated by covering existing EOS as well as the $\Gamma$-spectral posterior for GW170817 (2020arXiv200101747W). While this prior range nominally includes configurations with a NS maximum mass $M_{\rm max}$ smaller than the largest NS mass in our fiducial population ($m_{\rm max}=2.1M_{\odot}$), in practice when performing joint inference we always require consistency with the known massive NS and thus reject entirely any NS with maximum mass smaller than $m_{\rm max}$. [For the purposes of illustrating trends in single-constraint marginal likelihoods versus parameters, we at times relax this requirement and incorporate EOS realizations and marginal likelihoods which do not conform to this constraint.] Figure 2 shows our implied prior on $p(\rho)$ and nonrotating neutron star gravitational mass and radius. We use these two parametric approaches to qualitatively assess the kinds of systematic uncertainty exhibited among some of the commonly used parameterizations, such as piecewise polytropic (Read:2009iy), dynamical-piecewise polytropic (PhysRevD.106.103027; Carney:2018sdv), and similar spectral methods (Lindblom:2010bb; Lindblom:2018rfr; PhysRevD.105.063031) which adopt different basis functions or parameters (2023PhRvD.107l3017L).
@@
Bottom panel: Mass-Radius prior range for the EoSs in the above range. The grey shaded regions indicate various constraints on single-stable NS; ‘BH limit’ is where the Schwarzchild radius is larger than the NS radius, ‘Buchdahl limit’ is the range where an NS would need infinite pressure (PhysRev.116.1027), and ‘Causality limit’ is obtained by finding the maximally compact (M/R) solution for Mass-Radius in the limiting physical range of pressure-density (2012ARNPS..62..485L).
@@
We follow previous work that employs Gaussian quadrature to evaluate the energy-density at given pressure on the lalsuite implementation (LIGO-GW170817-SourceProperties; Carney:2018sdv). We adopt a fixed low-density equation of state with the first four terms in the expansion determining the EoS. This parameterization has been previously constrained by comparison of its input GW170817 (LIGO-GW170817-SourceProperties) and NICER measurements (Miller:2019cac; Miller:2021qha). For all EoSs, we solve the Tolman-Oppenheimer-Volkoff (TOV) equation (Oppenheimer:1939ne; Tolman:1939jz) for nonrotating neutron star structure in hydrostatic equilibrium using RePrimAnd (2024arXiv240411346K; wolfgang_kastaun_2023_7700296) to find one-parameter families of gravitational mass $M$, radius $R$, and tidal deformability $\Lambda$. The output for an equation of state $\alpha$ is characterized by a one-parameter family $R_{\alpha}(s),M_{\alpha}(s)$, where $s$ is a fiducial parameter characterizing our dimensionless prior. For simplicity, in this work we adopt the pseudo-enthalpy h, and only investigate the lowest stable branch of solutions.
2.8 Additional sources of external information
The nuclear EoS must be consistent with properties of nuclei (2012ARNPS..62..485L; 2021ARNPS..71..433L). Following (Miller:2021qha), one strategy for incorporating information about nuclear-density matter is to require the baryon density $n$ and energy density $\rho c^{2}$ to be consistent with constraints on the nuclear symmetry energy $S\simeq\rho/n-m_{n}c^{2}$, assumed to be known to be normally constrained with a mean of $32\penalty 10000\ {\rm MeV}$ and a standard deviation of $2\penalty 10000\ {\rm MeV}$ at saturation density, assumed to be 2.7 $\times 10^{14}\penalty 10000\ \rm{g\penalty 10000\ cm^{-3}}$ (PhysRevD.110.123009; PhysRevC.86.015803; 2019EPJA...55..117L). Figure 3 shows the marginal likelihood for different EoS realizations, plotted versus two EoS properties $R_{1.4}$ and $M_{\rm max}$. The symmetry energy constraint favors smaller radii at each fixed $M_{max}$.
Based on observations of AT2017gfo, several groups have argued that the observed electromagnetic energetics cannot be consistent with prompt collapse of the remnant to a black hole, using that qualitative constraint to infer an upper bound to the NS maximum mass; see, e.g., (2017ApJ...850L..34B; 2017LRR....20....3M; 2018ApJ...852L..25R; 2022PhRvD.105j3022K; 2018ApJ...852L..25R; 2023MNRAS.519.2615E) and references therein. Roughly speaking, these arguments suggest $M_{max}<2.3\penalty 10000\ {\rm M_{\odot}}$ (2018ApJ...852L..25R), extremely close to the mass inferred for J0952-0607 (2023MNRAS.519.2615E).
As noted in Section 2.3, the observation of AT2017gfo favors large amounts of ejecta. While these large ejecta masses can be accommodated with exceptional NS spins or mass ratios, these outcomes can also be produced with large neutron stars. Using this rationale, some groups have argued that the NS tidal deformability should be bounded below by approximately $\tilde{\Lambda}>400$ (2018ApJ...852L..29R). This proposed constraint on $\tilde{\Lambda}$ has comparable qualitative impact on the information obtained from GW170817 as our constraint that $q>0.9$ and $\chi_{i,z}>0$, as will be seen in Section 3.2 Figure 7. For brevity, we omit further discussion of this proposal.
A preliminary analysis with the inclusion of a recently measured galactic pulsar J0437 (2024ApJ...971L..20C) is discussed it in Appendix E.
2.9 Hyperparameter posterior pipeline: HyperPipe
To generate the posterior distribution from ${\cal L}_{\rm net}$ and $p({\cal Y})$ we introduce a generalization of the RIFT iterative inference strategy: HyperPipe. Starting with some proposed parameters ${\cal Y}_{\alpha}$, we evaluate ${\cal L}_{\rm net}({\cal Y}_{\alpha})\equiv{\cal L}_{net,\alpha}$; build an approximation $\hat{\cal L}({\cal Y})$; and use that approximation with an adaptive Monte Carlo integration program to generate both an estimated posterior distribution for $\cal Y$} as well as a new set of proposed parameters ${\cal Y}_{\alpha^{\prime}}$ meriting further investigation. This pipeline is able to adaptively explore the space of selected hyperparameters to resolve distinct features in the posterior distribution. The estimate is initiated in a small region of the hyperparameter space (in this case a subsection of the space of ${\cal Y}=\{(\gamma_{1},\gamma_{2}\ldots)\}$). We refer to the proposed parameter selection, ${\cal Y}_{\alpha}$, as the initial grid of parameters and the resultant physical properties as initial samples. Following the first iteration of evaluating likelihoods a posterior is estimated by performing Monte Carlo integration. Following this new parameters are sampled in the regions of the posterior with lower resolution in order to not exclude important finer details in addition to samples placed in exploratory regions. Following likelihood evaluations for these new points an updated posterior is evaluated by using prior and marginal likelihood information from samples of the previous iteration(s) and the current iteration, hence being better informed. Following this the process is repeated a few times where new sample points are generated and the posterior updated by Monte Carlo integration until the posterior does not deviate between iteration to iteration, resulting in a stable posterior estimate.